# Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense

### v4

@@ 1 Introduction
Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [39].” On the basis of that assumption, we designed the first effective white-box adaptive reversion attack, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible. Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense. Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier). If clean accuracy increases after reversion, the reversion succeeded. If it decreases after reversion (due to distortion), reversion failed.
@@ 2 Related Work
Lastly, Alfarra et al. [2] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer. However, the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
@@ 4 Adaptive Attack: Preemptive Reversion
A reversion attack is used to evaluate whether a preemptive defense is reversible. In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks. Preemptive Reversion is the only feasible approach to reverse preemptive defense in the white-box scenario. Nevertheless, without full knowledge of the backbone model and settings, our Fast Preemption defense is hard to reverse, so it is reliable.
@@ 5.1 Datasets, Classifier, and Backbone Models
We conducted experiments on two datasets: CIFAR-10 [24] and ImageNet [1]. For CIFAR-10, ResNeXt29-32x4d [19] and PreActResNet-18 [17] were selected as the classifier and backbone model, respectively. For ImageNet, DeiT Base [41] and ConvNeXt-L+ConvStem [40] were selected as the classifier and backbone model, respectively. The model architectures are listed in Tab. 3 of Appendix C.
@@ 6 Performance
In this section, we present the end-to-end evaluation results and comparisons with benchmarks. The results were obtained in the settings listed in Tab. 4 of Appendix D. The time cost was recorded with an NVIDIA A100 40GB GPU, as described in Appendix E.
@@ 8 Conclusion
We have presented an efficient transferable preemptive adversarial defense that we call Fast Preemption. We demonstrated that with its framework and Fore-Back cascade learning algorithm, Fast Preemption offers transferable protection against adversarial attacks among unknown victim models without requiring ground truths. In terms of the time cost, clean accuracy, and robust accuracy, Fast Preemption outperformed benchmark training-time, test-time, and preemptive defenses. We also demonstrated that the protection from Fast Preemption is irreversible unless the backbone model, algorithm, and settings of the defense are all compromised. This paper lays the theoretical foundation for a preemptive defense framework that can shape future research on each component’s optimization instead of end-to-end application in the relevant field.
Acknowledgments
This work was supported by JSPS KAKENHI Grants JP21H04907 and JP24H00732, by JST CREST Grants JPMJCR18A6 and JPMJCR20D3 including AIP challenge program, by JST AIP Acceleration Grant JPMJCR24U3, by JST K Program Grant JPMJKP24C2 Japan, and by the project for the development and demonstration of countermeasures against disinformation and misinformation on the Internet with the Ministry of Internal Affairs and Communications of Japan.
@@ Acknowledgments
This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix C), and the experimental configurations for both attacks and defenses (Appendix D). Additionally, it details the implementation specifics (Appendix E) and presents comprehensive ablation study results Appendix F, covering aspects such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
@@ Appendix A Related Work
Lastly, Alfarra et al. [2] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer. However, we believe that the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
@@
Appendix B Algorithms of Fast Preemption and
Preemptive Reversion
Appendix C Model Architectures
@@
Table 4: Experiment Settings.
        Type         Approach         Settings
        Adversarial         PGD         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$ or $L_{2}\ \epsilon=0.5^{1}/2^{2}$
        Attack         AutoAttack
                 Reversion         Attack         JPEG         Quality Factor = 50%
        DiffPure         Diffusion Timestep = $0.1^{1}/0.15^{2}$
        Fast Preemption (ours)         Same with our defense
        Defense         DiffPure         Diffusion Timestep = $0.1^{1}/0.15^{2}$
        Bi-Level         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
        $A^{5}$         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
  • 1 for CIFAR-10; 2 for ImageNet.
  • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
@@
Appendix D Experiment Settings
Appendix E Implementation
Appendix F Ablation Studies

### v5

@@ 1 Introduction
Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [43].” On the basis of that assumption, Bi-Level [29] first proposed a reversion attack but failed, claiming their defense was irreversible. However, our reversion attack successfully defeated their defense, establishing ours as the first effective adaptive reversion attack in the white-box setting, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible. Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense. Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier). If clean accuracy increases after reversion, the reversion succeeded. If it decreases after reversion (due to distortion), reversion failed.
@@ 2 Related Work
Lastly, Alfarra et al. [3] introduced a test-time defense mechanism incorporating preemptive defense as an anti-adversary layer guided by the output of a standard classifier. However, this defense can be circumvented if the classifier is compromised, a limitation explicitly acknowledged in their paper. This significant vulnerability undermines the practicality of using preemptive defense techniques in test-time scenarios. While they characterize this scenario as a “least realistic setting,” we contend that it aligns with the standard white-box attack paradigm, where adversaries have full access to the model. Under even more challenging assumptions, adversaries might also be aware of the defense mechanism itself (white-box reversion). Furthermore, the approach of Alfarra et al. [3] depends exclusively on forward propagation-based learning, which increases the risk of overfitting.
@@ 4 Adaptive Attack: Preemptive Reversion
A reversion attack is used to evaluate whether a preemptive defense is reversible. While Bi-Level [32] first proposed a reversion attack but failed, claiming their defense was irreversible, our reversion attack successfully defeated their defense, establishing ours as the first effective reversion attack. In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks. Preemptive Reversion is the only feasible approach to reverse preemptive defense in the white-box scenario. Nevertheless, without full knowledge of the backbone model and settings, our Fast Preemption defense is hard to reverse, so it is reliable.
@@ 5.1 Datasets, Classifier, and Backbone Models
We conducted experiments on two datasets: CIFAR-10 [26] and ImageNet [2]. For CIFAR-10, ResNeXt29-32x4d [21] and PreActResNet-18 [19] were selected as the classifier and backbone model, respectively. For ImageNet, DeiT Base [45] and ConvNeXt-L+ConvStem [44] were selected as the classifier and backbone model, respectively. The model architectures are listed in Tab. 4 of Appendix D.
@@ 6 Performance
In this section, we present the end-to-end evaluation results and comparisons with benchmarks. The results were obtained in the settings listed in Tab. 3 of Appendix C. The time cost was recorded with an NVIDIA A100 40GB GPU, as described in Appendix E.
@@ 8 Conclusion
We introduce an efficient and transferable preemptive adversarial defense, termed Fast Preemption. This approach establishes a novel layer in the defense hierarchy, Proactive Defense (ours) → Training-Time Defense → Test-Time Defense, distinct from conventional test-time and training-time defenses. The proposed mechanism is plug-and-play, transferable, and requires no prior knowledge, empowering users to safeguard media autonomously without reliance on third-party interventions. Moreover, our method addresses the efficiency limitations of proactive defenses, achieving processing speeds 700 times faster than existing approaches, a critical improvement for real-world applications where substantial delays (e.g., 28 seconds per image [32]) are infeasible. Finally, our approach maintains or enhances clean accuracy while delivering state-of-the-art robust accuracy, overcoming the common trade-off observed in other defense methods that degrade clean accuracy.
Nevertheless, our defense shows limitations against the proposed reversion attack in the white-box setting. While we design the defense allowing user customizing the backbone model (serving as a user-side key) to counter reversion attacks, simplifying user requirements is a priority for future work. Further research should also explore the relationship between backbone models and transferability, as discussed in Sec. F.2.
@@
This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix D), and the experimental configurations for both attacks and defenses (Appendix C). Additionally, it details the implementation specifics (Appendix E) and presents comprehensive ablation study results Appendix F, covering aspects such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
@@ Appendix A Related Work
Lastly, Alfarra et al. [3] introduced a test-time defense mechanism incorporating preemptive defense as an anti-adversary layer guided by the output of a standard classifier. However, this defense can be circumvented if the classifier is compromised, a limitation explicitly acknowledged in their paper. This significant vulnerability undermines the practicality of using preemptive defense techniques in test-time scenarios. While they characterize this scenario as a “least realistic setting,” we contend that it aligns with the standard white-box attack paradigm, where adversaries have full access to the model. Under even more challenging assumptions, adversaries might also be aware of the defense mechanism itself (white-box reversion). Furthermore, the approach of Alfarra et al. [3] depends exclusively on forward propagation-based learning, which increases the risk of overfitting.
Appendix B Algorithms of Fast Preemption and
Preemptive Reversion
@@ Appendix B Algorithms of Fast Preemption and
Table 3: Experiment Settings.
       Type        Approach        Settings
               Adversarial        Attack        PGD [31]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$ or $L_{2}\ \epsilon=0.5^{1}/2^{2}$
       AutoAttack [11]
       CW [7]
       DeCoW [28]
       MCG [53]        Maximum Queries = 20,000
               Reversion        Attack        JPEG        Quality Factor = 50%
       DiffPure [34]        Diffusion Timestep = $0.1^{1}/0.15^{2}$
       Fast Preemption (ours)        Same with our defense
       Defense        DiffPure [34]        Diffusion Timestep = $0.1^{1}/0.15^{2}$
       Bi-Level [32]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
       $A^{5}$ [17]        $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
  • 1 for CIFAR-10; 2 for ImageNet.
  • 3 $\eta$ must meet $\eta\cdot(T_{\rightarrow}+T_{\leftarrow})>1$.
Appendix C Experiment Settings
Appendix D Model Architectures
@@
Appendix E Implementation
Appendix F Ablation Studies
@@ F.4 Optimization Algorithm
Appendix G Evaluation across More Models
In the main paper, we concentrated on comparisons with top-performing adversarial training methods [6, 29]. In this section, we extend our evaluation to include defenses tested across nine additional victim models on CIFAR-10 [26] from RobustBench [12] in a transferable context. These models encompass diverse architectures representative of platforms such as Twitter and Facebook, differing from both the classifier [21] and the backbone model [19].
As shown in Tab. 8, our method demonstrates significant performance gains compared to no proactive defense and the SOTA proactive defense [32]. Specifically, we achieve average improvements of 4.2% and 4.3% in clean accuracy, 17.8% and 5% in robust accuracy, and a 700-fold increase in speed, respectively. Furthermore, compared using the SOTA adversarial training model [6] as the backbone model, our chosen [19] yields an average increase of 0.7% in clean accuracy, 2.5% in robust accuracy, and a tenfold improvement in processing speed.
Table 8: Transferable clean and robust accuracy (%) across various victim models against AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$. All proactive defenses were conducted using classifier and backbone models distinct from the victim models, even when the architectures were similar.
Model Metric No Proactive Defense Bi-Level [32] Fast Preemption (ours)
Defense Learning Backbone N/A Bi-Level [32] WideResNet-94-16 [6] PreActResNet-18 [19]
Time Cost (s) N/A 28 0.35 0.04
RaWideResNet-70-16 [35] Clean 93.5 92.1 95.5 95.4
Robust 71.4 81.3 86.8 85.9
WideResNet-70-16 [50] Clean 93.3 91.5 95.5 95.3
Robust 71.2 80.9 86.4 86.0
WideResNet-34-10 [36] Clean 91.5 90.2 95.1 95.1
Robust 62.1 76.2 78.7 80.8
WideResNet-28-10 [52] Clean 93.4 91.5 95.4 95.6
Robust 64.8 77.7 81.3 83.2
PreActResNet-18 [51] Clean 84.2 87.3 90.1 91.6
Robust 42.4 54.7 55.4 61.2
ResNest152 [41] Clean 86.9 87.6 91.2 92.4
Robust 62.4 73.2 75.8 78.2
ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
Robust 52.0 68.1 66.7 72.0
ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
Robust 51.4 67.5 67.8 72.8
XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
Robust 57.9 71.3 74.6 76.0
Average Clean 89.9 89.8 93.4 94.1
Robust 59.5 72.3 74.8 77.3
  • Bold indicates the best performance.
Appendix H Defense against More Attacks
Our primary evaluation adheres to the RobustBench [12] protocol, employing the strongest attack, AutoAttack [11]. As detailed in Sec. 5.2, this includes tests with standard and randomized AutoAttack variants, encompassing white-box APGD [11], FAB [10], and black-box Square [4] attacks, alongside EOT settings [5].
In this section, we extend the evaluation to include three additional attacks: white-box Carlini-Wagner (CW) [7], black-box Deformation-Constrained Warping (DeCoW) [28], and black-box Meta Conditional Generator (MCG) [53], demonstrating the broad applicability and generalizability of our defense against diverse attack strategies.
As shown in Tab. 9, our defense achieves robust accuracy of 90.3%, 93.6%, and 87.3% against CW, DeCoW, and MCG attacks, respectively, outperforming SOTA adversarial training [6] and proactive defenses [32]. Notably, these accuracies exceed those obtained against AutoAttack, further highlighting the robustness and reliability of our primary evaluation.
Table 9: Robust accuracy (%) against various adversarial attacks with $L_{\infty}$-norm perturbations and $\epsilon=8/255$.
      Threat Model       Attack       WideResNet-94-16 [6]       Bi-Level [32]       Fast Preemption (ours)
      White-Box       AutoAttack (Standard) [11]       74.0       82.1       86.3
      CW [7]       80.8       85.4       90.3
      Black-Box       DeCoW [28]       88.8       89.7       93.6
      MCG [53]       76.1       83.3       87.3
  • Bold indicates the best performance.
Appendix I Visual Quality
The noise budgets utilized (as detailed in Tab. 3) adhere to standard practices in adversarial machine learning, ensuring the preservation of visual quality and minimizing the likelihood of detection. As illustrated in Fig. 10, for 1,000 images from CIFAR-10 and ImageNet, Structural Similarity Index Measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] scores confirm that the visual quality of our robust examples (REs) is sufficiently high to serve as effective substitutes for the original images. Furthermore, our REs exhibit comparable (often superior) visual quality relative to adversarial examples (AEs) and outputs of existing defense methods.
Figure 10: Visual quality of robust examples (REs) generated by our Fast Preemptive defense, compared with adversarial examples (AEs) from AutoAttack and REs from the SOTA proactive defense [32]. All samples adhere to the same $L_{\infty}$-norm constraint ($\epsilon=8/255$) relative to the original images. SSIM and LPIPS scores demonstrate that the visual quality of our REs is sufficiently high to serve as replacements for the original images, and is comparable to (and often slightly better than) those of the adversarial attack and existing defense methods.
I.1 Structural Similarity Index (SSIM)
The SSIM is a metric for measuring similarity between two images. It considers luminance, contrast, and structural information. The SSIM between two images $x$ (reference) and $y$ (distorted) is defined as:
$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (15)
where:
  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
  • $\sigma_{xy}$: Covariance between $x$ and $y$.
  • $C_{1}$ and $C_{2}$: Small constants added to stabilize the metric in the presence of weak denominators, defined as:
    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
    where $L$ is the dynamic range of the pixel values (e.g., 255 for 8-bit images), and $K_{1}$ and $K_{2}$ are small positive constants (commonly, $K_{1}=0.01$ and $K_{2}=0.03$).
The SSIM value ranges from -1 to 1, where:
  • 1: Perfect similarity.
  • 0: No similarity.
  • Negative values indicate dissimilarity.
SSIM is typically calculated locally using a sliding window across the image, and the overall SSIM is the average of local SSIM values. Higher SSIM values denote subtler deviations from the reference images.
I.2 Learned Perceptual Image Patch Similarity (LPIPS)
The LPIPS is a perceptual similarity measure that evaluates the perceptual distance between two images. It uses deep neural network features to compare the visual content and correlates better with human perception than traditional metrics like SSIM.
The LPIPS distance between two images $x$ (reference) and $y$ (distorted) is computed as:
$LPIPS(x,y)=$ (17)
$\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}w_{l}\|\hat{\phi}_{l}(x)_{h,w}-\hat{\phi}_{l}(y)_{h,w}\|_{2}^{2},$
where:
  • $l$: Layer index in the deep network.
  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
  • $w_{l}$: Learned scalar weights for each layer $l$ (trained to match human judgments).
LPIPS uses a pretrained deep neural network to extract feature maps from intermediate layers. The metric normalizes the feature maps and computes the difference between them, weighted according to the learned weights $w_{l}$. The LPIPS value typically ranges from 00 (perfect perceptual similarity) to higher values indicating greater dissimilarity. LPIPS provides a perceptually grounded comparison, making it especially useful for evaluating subtle changes in visual quality.